# Planet-Planet Scattering and ZLK Migration – The Dynamical History of HAT-P-11

### v1

@@ 1 Introduction
The mid-K dwarf HAT-P-11 hosts an intriguing system that has been the subject of much interest in recent years. The system has two known planets. The first, HAT-P-11 b, is a close-in ($a_{\mathrm{b}}=0.525$ AU) eccentric ($e_{\mathrm{b}}=0.218$) super-Neptune ($m_{\mathrm{b}}=23.4\>\mathrm{M}_{E}$) first identified by Bakos et al. (2010). Follow-up Rossiter-McLaughlin (Rossiter, 1924; McLaughlin, 1924) analysis of the system by Hirano et al. (2011) and Sanchis-Ojeda & Winn (2011) revealed that the orbit of HAT-P-11 b is polar ($\psi_{\mathrm{Ab}}=106^{\circ}$), which makes it a member of the potential population of perpendicular planets (Albrecht et al., 2021; Dong & Foreman-Mackey, 2023; Siegel et al., 2023).
HAT-P-11 b presents an interesting dynamical puzzle. Planets as close-in as HAT-P-11 b are expected to be in circular orbits due to tidal forces. The fact that HAT-P-11 b is significantly eccentric, coupled with its unusual perpendicular orbit, points to a dynamically hot history. Yee et al. (2018) significantly advanced our understanding of the dynamical history of the system with their discovery of the second planet in the system, HAT-P-11 c, an eccentric ($e_{\mathrm{c}}\approx 0.6$) super-Jupiter ($m_{\mathrm{c}}\sin i\approx 1.6\>\mathrm{M}_{J}$) on an $a_{\mathrm{c}}\approx 4.1$ AU orbit. Their dynamical analysis of the system concluded that with this additional companion, both the eccentricity and stellar obliquity of HAT-P-11 b could be explained on the condition that HAT-P-11 c was also misaligned. At the time, the orbital inclination of HAT-P-11 c was unconstrained.
The final piece of the puzzle was provided by An et. al (in prep, hereafter Paper 1), who combined radial velocities (RVs), astrometric data from Hipparcos and Gaia, and projected obliquity constraints to confirm that the orbits of HAT-P-11 b and HAT-P-11 c are indeed misaligned, with a bimodal inclination posterior for HAT-P-11 c at values of $i_{\mathrm{c}}\approx 34^{\circ}$ or $i_{\mathrm{c}}\approx 144^{\circ}$. The mutual inclination between the orbits of HAT-P-11 c and HAT-P-11 b, and the stellar obliquity of HAT-P-11 c, were both constrained to be $>$30 at $3\sigma$-equivalent significance, i.e., with significant misalignments measured for both. Given the constraints provided by Paper 1, all relevant orbital quantities to the dynamics of the system are well-constrained and are presented in Table 1. A schematic view of the orbits is shown in Figure 1. We are now in a position to postulate the complete dynamical history of HAT-P-11.
@@
HAT-P-11 baaHirano et al. (2011) and Huber et al. (2017) give different values for stellar obliquity and eccentricity, respectively.
@@
Mass (MJ) $3.06^{+0.42}_{-0.42}$ Paper 1
a (AU) $4.192^{+0.07}_{-0.07}$ Paper 1
e $0.56^{+0.036}_{-0.035}$ Paper 1
$\psi_{\mathrm{Ac}}(^{\circ})$ 43 to 138bbRange enclosing 95% of the (highly non-Gaussian) posterior Paper 1
$\psi_{\mathrm{bc}}(^{\circ})$ 50 to 130bbRange enclosing 95% of the (highly non-Gaussian) posterior Paper 1
@@
Both planets in the HAT-P-11 system represent dynamical puzzles, and the dynamical history of the system has been explored by Petrovich et al. (2020); Pu & Lai (2021) among others. In this work we propose a two-step formation mechanism: we posit that both HAT-P-11 b and HAT-P-11 c initially formed in circular aligned orbits, along with two other large planets. Through dynamical instability and subsequent planet-planet scattering, both additional planets were ejected from the system. As a result, HAT-P-11 b and HAT-P-11 c were left on eccentric and misaligned orbits, which triggered von-Zeipel-Lidov-Kozai (ZLK) migration in HAT-P-11 b. Through these mechanisms, all aspects of the present-day configuration of the system are reproduced.
@@ 2 Planet-Planet Scattering
The high eccentricity of HAT-P-11 c, coupled with the misaligned stellar obliquity reported in Paper 1, is the first dynamical puzzle. Naively, one would expect planets to form in circular, aligned orbits, thanks to strong eccentricity damping within the protoplanetary disk. Any deviation from this paradigm is believed to be the imprint of a dynamically active history. A number of well-understood mechanisms are known to excite the eccentricities and inclinations of planets, including but not limited to the von-Zeipel-Lidov-Kozai (ZLK) effect (von Zeipel, 1910; Lidov, 1962; Kozai, 1962) and secular chaos (Wu & Lithwick, 2011). Both of these mechanisms require an additional undetected companion in the system, something for which there is no direct evidence. The existing radial velocity data are satisfactorily explained by two planets with no residual trend, while the measured astrometric acceleration is similarly consistent with the two known planets (Paper 1). To evade these constraints, an additional planet would have to be low-mass and/or very widely separated and hence be dynamically insignificant. Planet-disk interactions during planetary migration can excite eccentricities without the need for a companion, but this mechanism has been shown to be limited to relatively low eccentricities, of order $e\sim 0.1$ (Duffell & Chiang, 2015), and it also cannot generate stellar obliquity misalignment.
In the absence of a more distant perturber, the most promising pathway is planet-planet scattering, where close encounters between pairs of planets lead to strong gravitational interactions that lead to large eccentricities and obliquities. While additional planets would be required for this mechanism, violent scattering is prone to ejecting planets or scattering them to wide orbits where they would be difficult to detect (e.g., Chatterjee et al., 2008; Carrera et al., 2019; Frelikh et al., 2019). In this section, we explore the generation of HAT-P-11 c’s significant eccentricity and orbital misalignment through N-body simulations of planet-planet scattering.
@@ 2.1 Theoretical Background
The stability of systems involving only two planets is well-understood analytically (Gladman, 1993). Studies such as Ford et al. (2001) and Petrovich et al. (2014) have shown that two-planet systems can attain a wide range of eccentricities but cannot reach high mutual inclinations. Three-body scattering, on the other hand, can both excite eccentricity to arbitrarily large values and raise inclination up to $90^{\circ}$ (Chatterjee et al., 2008; Carrera et al., 2019). In the stellar obliquity posterior derived in Paper 1, the $95\%$ confidence interval is given by 43°43°43  ° to 138°138°138  °; thus the $\sim$40 lower limit requires three-body scattering to be invoked.
The stability of three-body systems cannot be obtained analytically. From numerical studies, system stability is both highly chaotic and dependent on mean-motion resonances (Marzari, 2014; Rath et al., 2022, e.g.,). However, to first-order the time to dynamical instability grows logarithmically with mutual separation (Marzari & Weidenschilling, 2002). The mutual separation is commonly parameterized by the mutual Hill radius:
@@
and $\Delta$ denotes the separation in unit of the mutual Hill radius,
@@
The stability timescale grows as the separation between the planets increases (e.g., Chambers et al., 1996; Pu & Wu, 2015; Tamayo et al., 2020).
@@ 2.2.1 Three-Planet Simulations
We perform 1000 three-planet simulations using the IAS15 integrator (Rein & Spiegel, 2014) in the REBOUND N-body integrator package (Rein & Liu, 2012). We use the adaptive timestep criterion described in Pham et al. (2024). The setup of our simulations is as follows: we consider three identical planets with $m_{1}=m_{2}=m_{3}=3.06\>M_{\mathrm{J}}$, the best-fit value for HAT-P-11 c’s mass. Three-body scattering tends to result in a single surviving planet, or additional survivors on highly eccentric/marginally unbound orbits (e.g Carrera et al., 2019). In both cases, the orbital energy of the non-innermost planets is nearly zero. Conservation of energy thus demands
@@
We set $a_{f}=4.192$ AU, the present-day semi-major axis of HAT-P-11 c. In the interest of minimizing computation time, we set $\Delta=3$ to induce rapid instability and scattering in our systems. This is unrealistically compact for primordial planetary systems, where separations would be expected to be around 5-10 mutual Hill radii (Raymond & Morbidelli, 2022). Changing the mutual separation of the planets does not qualitatively affect the scattering outcome, only the timescale until dynamical instability (Chatterjee et al., 2008). The above criterion hence informs the initial semi-major axes of the three planets: $a_{1,i}=8.838$ AU, $a_{2,i}=13.281$ AU, $a_{3,i}=19.959$ AU. In each simulation, the other relevant orbital parameters for each planet are chosen at random from the following uniform distributions:
@@
Note that we do not track the motion of the stellar spin axis in our simulations. Rather, we assume that the star’s initial spin axis points along the $z$-axis and does not evolve significantly over the course of our simulations. With this assumption, the orbital inclination $i$ can be used as a proxy for the stellar obliquity $\psi_{\mathrm{Ac}}$. We account for collisions via merging planets (conserving mass, volume and momentum but not energy) using the reb_collision_resolve_merge collision module. We also remove particles if they exceed a distance of $10^{4}$ AU from the origin, using the exit_max_distance feature.
Figure 2 shows an example of one of our simulations. The system rapidly enters a regime of dynamical instability, as expected from the initialization of the planets at small mutual Hill radii. This epoch is characterized by sudden perturbations in semimajor axis, eccentricity and inclination triggered by close encounters. At around $5$ Myr, the first planet is completely ejected from the system. At this point, both remaining planets in the system gain significant eccentricities and inclinations. The system remains stable over the remaining duration of the simulation. The final orbital elements of HAT-P-11 c are $a_{\mathrm{c}}=4.24$ AU, $e_{\rm c}=0.58$, and $\psi_{\mathrm{Ac}}=68.5^{\circ}$, while the orbital elements of the surviving outer planet are $a_{\mathrm{outer}}=240$ AU, $e_{\mathrm{outer}}=0.73$, and $\psi_{\mathrm{outer}}=$$$ 26.3  °.
Figure 2: Example three-body scattering results for the origin of HAT-P-11 c. The thick maroon line and the two thin blue lines represent the orbital elements of HAT-P-11 c and the two ejected planets (EP), respectively. The subplots show the semimajor axis, eccentricity, and stellar obliquity evolution of the three bodies, from top to bottom. When a planet is fully ejected from the system (heliocentric distance $>10^{4}$ AU) or becomes unbound $(a<0)$, the line associated with the evolution of that planet is terminated. The black dot on the right-hand side of each subplot denotes our derived values for the present-day system, with the error bar representing the $1\sigma$ posteriors for semimajor axis and eccentricity and the $2\sigma$ posterior for stellar obliquity.
In Figure 3 we report statistics from the entire ensemble of simulations. We plot histograms of the distributions in $a_{\mathrm{c}}$, $e_{\rm c}$ and $\psi_{\rm Ac}$ of the innermost surviving planet in each simulation, as well as the present-day $1\sigma$ constraints ($2\sigma$ for $\psi_{\rm Ac}$) on the orbit of HAT-P-11 c from Paper 1. The simulation results are as expected: $a_{\mathrm{c}}$ is strongly peaked near the present-day value, and both $e_{\rm c}$ and $i_{\mathrm{c}}$ show a wide range of values that easily encompass the constraints. This shows that our system setup is capable of consistently producing orbits similar to that of HAT-P-11 c.
@@
Of the $1000$ simulations we ran, the number with $1$, $2$ and $3$ surviving planets were $100,878$ and $22$, respectively. The most common outcome by far is a system with two surviving planets, which would imply an external undiscovered planet in the HAT-P-11 system. This is not in tension with observations, because the extraneous surviving planets are generally very far out, and hence they would both be undetectable and play a negligible role in the future dynamics of the system. In our simulations there are a total of $908$ extraneous planets that survive. The $1\sigma$ semimajor axis distribution of these planets is $a_{\mathrm{extra}}=80^{+150}_{-37}$ AU, while the perihelion distribution of these $a_{p,\mathrm{extra}}=32^{+20}_{-12}$ AU. There is only one simulation where there is an orbit crossing between any of the extra planets and the surviving HAT-P-11 c. We do not analyze the stability of the surviving systems in-depth; an analytic criteria such as the one introduced by Hadden & Lithwick (2018) could be used to this end. We conclude that planet-planet scattering in a three planet system well-reproduces the unusual stellar obliquity and eccentricity of HAT-P-11 c.

### v2

@@ 1 Introduction
The mid-K dwarf HAT-P-11 hosts an intriguing system that has been the subject of much interest in recent years. The system has two known planets. The first, HAT-P-11 b, is a close-in ($a_{\mathrm{b}}=0.0525$ AU) eccentric ($e_{\mathrm{b}}=0.218$) super-Neptune ($m_{\mathrm{b}}=23.4\>\mathrm{M}_{E}$) first identified by Bakos et al. (2010). Follow-up Rossiter-McLaughlin (Rossiter, 1924; McLaughlin, 1924) analysis of the system by Hirano et al. (2011) and Sanchis-Ojeda & Winn (2011) revealed that the orbit of HAT-P-11 b is polar ($\psi_{\mathrm{Ab}}=106^{\circ}$), which makes it a member of the potential population of perpendicular planets (Albrecht et al., 2021; Dong & Foreman-Mackey, 2023; Siegel et al., 2023).
HAT-P-11 b presents an interesting dynamical puzzle. Planets as close-in as HAT-P-11 b are expected to be on circular orbits due to tidal forces. The fact that HAT-P-11 b is significantly eccentric, coupled with its unusual perpendicular orbit, points to a dynamically hot history. Yee et al. (2018) greatly advanced our understanding of the dynamical history of the system with their discovery of the second planet in the system, HAT-P-11 c, an eccentric ($e_{\mathrm{c}}\approx 0.6$) super-Jupiter ($m_{\mathrm{c}}\sin i\approx 1.6\>\mathrm{M}_{J}$) on an $a_{\mathrm{c}}\approx 4.1$ AU orbit. Their dynamical analysis of the system concluded that with this additional companion, both the eccentricity and spin-orbit misalignment of HAT-P-11 b could be explained on the condition that HAT-P-11 c was also misaligned. At the time, the orbital inclination of HAT-P-11 c was unconstrained.
The final piece of puzzle, the misalignment between the orbits of HAT-P-11 b and HAT-P-11 c, was initially explored by Xuan & Wyatt (2020), in which they derived a bimodal inclination distribution for HAT-P-11 c, providing evidence of a large mutual inclination between planet b and c. An et al. (2024) (hereafter Paper 1) combined astrometry data from Hipparcos and Gaia with radial velocity data from Keck/HIRES to constrain the orbit of HAT-P-11 c, including a more precise bimodal constraint on inclination. All relevant orbital quantities of the system are well-constrained and are presented in Table 1. In Paper 1, the two inclination modes are treated separately, together with the projected stellar spin constrained by Sanchis-Ojeda & Winn (2011), to calculate the alignment of all three main angular momentum vectors in the HAT-P-11 system. Paper 1 reports the first measurement of the spin-orbit misalignment of HAT-P-11 c, and significantly improved measurements of the mutual inclination between HAT-P-11 b and HAT-P-11 c. A schematic view of the orbits is shown in Figure 1. We are now in a position to postulate the complete dynamical history of HAT-P-11.
@@
HAT-P-11 baaHirano et al. (2011) and Huber et al. (2017) give slightly different values for spin-orbit misalignment and eccentricity, respectively.
@@
Mass (MJ) ${2.68\pm 0.41}$ Paper 1
a (AU) $4.10\pm 0.06$ Paper 1
e $0.652\pm 0.017$ Paper 1
$\psi_{\mathrm{Ac}}(^{\circ})$ 45 to 138bbRange enclosing 95% of the (highly non-Gaussian) posterior Paper 1
$\psi_{\mathrm{bc}}(^{\circ})$ 49 to 131bbRange enclosing 95% of the (highly non-Gaussian) posterior Paper 1
@@
Both planets in the HAT-P-11 system represent dynamical puzzles, and the formation history of the system has been explored by Petrovich et al. (2020); Pu & Lai (2021) among others. In this work we propose a two-step formation mechanism: we posit that both HAT-P-11 b and HAT-P-11 c initially formed in circular aligned orbits, along with two other large planets. Through dynamical instability and subsequent planet-planet scattering, both additional planets were ejected from the system. As a result, HAT-P-11 b and HAT-P-11 c were left on eccentric and misaligned orbits, which triggered von-Zeipel-Lidov-Kozai (ZLK) migration in HAT-P-11 b. Through these mechanisms, all aspects of the present-day configuration of the system are reproduced.
@@ 2 Planet-Planet Scattering
The high eccentricity of HAT-P-11 c, coupled with the significant spin-orbit misalignment reported in Paper 1, is the first dynamical puzzle. Naively, one would expect planets to form in circular, aligned orbits, thanks to strong eccentricity damping within the protoplanetary disk. Any deviation from this paradigm is believed to be the imprint of a dynamically active history. A number of well-understood mechanisms are known to excite the eccentricities and inclinations of planets, including but not limited to the von-Zeipel-Lidov-Kozai (ZLK) effect (von Zeipel, 1910; Lidov, 1962; Kozai, 1962) and secular chaos (Wu & Lithwick, 2011). Both of these mechanisms require an additional undetected companion in the system, something for which there is no direct evidence in the HAT-P-11 system. The existing radial velocity data are satisfactorily explained by two planets with no residual trend, and the measured astrometric acceleration is similarly consistent with the two known planets (Paper 1). To evade these constraints, an additional planet would have to be low-mass and/or very widely separated and hence be dynamically insignificant. Planet-disk interactions during planetary migration can excite eccentricities without the need for a companion, but this mechanism cannot generate the spin-orbit misalignment.
In the absence of a more distant perturber, the most promising pathway is planet-planet scattering, where close encounters between pairs of planets lead to strong gravitational interactions that produce large eccentricities and obliquities. While additional planets would be required for this mechanism, violent scattering is prone to ejecting planets or scattering them to wide orbits where they would be difficult to detect (e.g., Chatterjee et al., 2008; Carrera et al., 2019; Frelikh et al., 2019). In this section, we explore the generation of HAT-P-11 c’s significant eccentricity and orbital misalignment through N-body simulations of planet-planet scattering.
@@ 2.1 Theoretical Background
The stability of systems involving only two planets is well-understood analytically (Gladman, 1993). Studies such as Ford et al. (2001), Petrovich et al. (2014) and Gratia & Fabrycky (2017) have shown that two-planet systems can attain a wide range of eccentricities but very rarely high mutual inclinations. Three-body scattering, on the other hand, can both excite eccentricity to arbitrarily large values and raise inclination up to $90^{\circ}$ (Chatterjee et al., 2008; Jurić & Tremaine, 2008; Carrera et al., 2019). In the spin-orbit misalignment posterior derived in Paper 1, the $95\%$ confidence interval is given by 45°45°45  ° to 138°138°138  °; thus the $\sim$40 lower limit requires three-body scattering to be invoked.
The stability of three-body systems cannot be obtained analytically. From numerical studies, system stability is both highly chaotic and dependent on mean-motion resonances (e.g., Marzari, 2014; Rath et al., 2022). However, to first-order the time to dynamical instability grows logarithmically with mutual separation (Marzari & Weidenschilling, 2002). The mutual separation is commonly parameterized by the mutual Hill radius (though note that other criteria often offer better predictions, e.g Lammers et al. (2024)):
@@
where $a_{i}$, $m_{i}$ are the semimajor axes and masses of the $i$th planet and $M_{*}$ is the mass of the star. The $\Delta$ parameter denotes the separation in units of the mutual Hill radius,
@@
The instability timescale grows as the separation between the planets increases (e.g., Chambers et al., 1996; Pu & Wu, 2015; Tamayo et al., 2020).
@@ 2.2.1 Three-Planet Simulations
We perform 1000 three-planet simulations using the IAS15 integrator (Rein & Spiegel, 2014) in the REBOUND N-body integrator package (Rein & Liu, 2012). We use the adaptive timestep criterion described in Pham et al. (2024). The setup of our simulations is as follows: we consider three identical planets with $m_{1}=m_{2}=m_{3}=2.68\>M_{\mathrm{J}}$, the best-fit value for HAT-P-11 c’s mass. We note that our results are not sensitive to the masses of the three planets. We also ran a set of simulations with $m_{1}=m_{\mathrm{c}},m_{2}=m_{\mathrm{c}}/2$ and $m_{3}=m_{\mathrm{c}}/4$ following the prescription of Anderson et al. (2020), and found no significant differences at the population level.
Three-body scattering tends to result in a single surviving planet, or additional survivors on highly eccentric/marginally unbound orbits (e.g Carrera et al., 2019). In both cases, the orbital energy of the non-innermost planets is nearly zero. Conservation of energy thus demands
@@
We set $a_{f}=4.10$ AU, the present-day semi-major axis of HAT-P-11 c. In the interest of minimizing computation time, we set $\Delta=3$ to induce rapid instability and scattering in our systems. This is unrealistically compact for primordial planetary systems, where separations would be expected to be around 5-10 mutual Hill radii (Raymond & Morbidelli, 2022). Changing the mutual separation of the planets does not qualitatively affect the scattering outcome, only the timescale until dynamical instability (e.g Chatterjee et al., 2008; Anderson et al., 2020). The above criterion hence informs the initial semi-major axes of the three planets: $a_{1,i}=8.759$ AU, $a_{2,i}=12.930$ AU, $a_{3,i}=19.086$ AU. In each simulation, the other relevant orbital parameters for each planet are chosen at random from the following uniform distributions:
@@
Note that we do not track the motion of the stellar spin axis in our simulations. Rather, we assume that the star’s initial spin axis points along the $z$-axis and does not evolve significantly over the course of our simulations. With this assumption, the orbital inclination $i$ can be used as a proxy for the spin-orbit misalignment $\psi_{\mathrm{Ac}}$. We account for collisions via merging planets (conserving mass, volume and momentum but not energy) using the reb_collision_resolve_merge collision module. We also remove particles if they exceed a distance of $10^{3}$ AU from the origin, using the exit_max_distance feature.
Figure 2 shows an example of one of our simulations. The system rapidly enters a regime of dynamical instability, as expected from the initialization of the planets at small mutual Hill radii. This epoch is characterized by sudden perturbations in semimajor axis, eccentricity and inclination triggered by close encounters. Within 1 Myr, one planet is completed ejected from the system. The remaining two planets continue to experience perturbations in eccentricity and inclination, until a second planet is ejected from the system around $5$ Myr. At this point, HAT-P-11 c has gained significant eccentricity and inclination. The final orbital elements of HAT-P-11 c are $a_{\mathrm{c}}=4.12$ AU, $e_{\rm c}=0.639$, and $\psi_{\mathrm{Ac}}=65.3^{\circ}$, which are all time-averaged values over the last $5$ Myr of the simulation.
Figure 2: Example three-body scattering results for the origin of HAT-P-11 c. The thick maroon line and the two thin blue lines represent the orbital elements of HAT-P-11 c and the two ejected planets (EP), respectively. The subplots show the semimajor axis, eccentricity, and spin-orbit misalignment evolution of the three bodies, from top to bottom. When a planet is fully ejected from the system (heliocentric distance $>10^{3}$ AU) or becomes unbound $(a<0)$, the line associated with the evolution of that planet is terminated. The black dot on the right-hand side of each subplot denotes our derived values for the present-day system, with the error bar representing the $1\sigma$ posteriors for semimajor axis and eccentricity and the $2\sigma$ posterior for spin-orbit misalignment.
In Figure 3 we report statistics from the entire ensemble of simulations. We plot histograms of the distributions in $a_{\mathrm{c}}$, $e_{\rm c}$ and $\psi_{\rm Ac}$ of the innermost surviving planet in each simulation, as well as the present-day $1\sigma$ constraints ($2\sigma$ for $\psi_{\rm Ac}$) on the orbit of HAT-P-11 c from Paper 1. The simulation results are as expected: $a_{\mathrm{c}}$ is strongly peaked near the present-day value, and both $e_{\rm c}$ and $\psi_{\mathrm{Ac}}$ show a wide range of values that easily encompass the constraints. This shows that our system setup is capable of consistently producing orbits similar to that of HAT-P-11 c.
@@
Of the $1000$ simulations we ran, the number with $1$, $2$ and $3$ surviving planets were $197,802$ and $1$, respectively. The most common outcome by far is a system with two surviving planets, which would imply an external undiscovered planet in the HAT-P-11 system.
The extra planet is not in tension with observations, because the extraneous surviving planet is generally very far out and hence would be undetectable. In some cases the extra planet produces long-term secular oscillations in the orbital elements of HAT-P-11 c, but this does not impact the system’s stability. In our simulations there are a total of $908$ extraneous planets that survive. The $1\sigma$ semimajor axis distribution of these planets is $a_{\mathrm{extra}}=72^{+123}_{-37}$ AU, while the perihelion distribution of these $a_{p,\mathrm{extra}}=28^{+17}_{-9}$ AU. There are only 5 simulations where there is an orbit crossing between any of the extra planets and the surviving HAT-P-11 c. We do not analyze the stability of the surviving systems in-depth; an analytic criterion such as the one introduced by Hadden & Lithwick (2018) could be used to this end. We conclude that planet-planet scattering in a three planet system well-reproduces the unusual spin-orbit misalignment and eccentricity of HAT-P-11 c.